OpenStudy (ayonnaleflore):

Factor. 12x2y – 3x2y2 – 15xy 3(12x2y – xy – 15xy) 3(4xy – x2y – 15xy) 3xy(4x – xy – 5) 3x2y2(4xy – xy – 5)

OpenStudy (dayakar):

do u have any idea

OpenStudy (ayonnaleflore):

none at all

OpenStudy (dayakar):

12x^2y = 2*2*3 *x*x*y [written product of prime factors of a first term] can u write another two terms like this

OpenStudy (ayonnaleflore):

3x^2y^2 = 2*x*x*2*y*y

OpenStudy (ayonnaleflore):

15xy = 5*3*x*y

OpenStudy (dayakar):



OpenStudy (dayakar):

is it clear

OpenStudy (ayonnaleflore):

so my answer would be 3xy(4x – xy – 5)

OpenStudy (dayakar):

yes

OpenStudy (ayonnaleflore):

ok thank you